# Loss Shaping Constraints for Long-Term Time Series Forecasting

### v1

@@ 5 Experiments
In order to explore the properties of loss shaping, we turn to the open source benchmark datasets Electricity Consuming Load (ECL) [^2] ECL dataset was acquired at https://archive.ics.uci.edu/ml/ datasets/ElectricityLoadDiagrams20112014, Weather  [^3] Weather dataset was acquired at https://www.ncei.noaa.gov/ data/local-climatological-data/ and Exchange (Lai et al., 2018) to train three popular transformer architectures for time series prediction, namely, Reformer (Kitaev et al., 2019) and Autoformer (Wu et al., 2021), Informer (Zhou et al., 2021) and a vanilla transformer. We contrast their behavior in constrained and (standard) unconstrained learning settings. We follow the same setup and hyperparameters used in (Kitaev et al., 2019; Wu et al., 2021; Zhou et al., 2021). Data is split into train, validation, and test chronologically with a ratio of 7:1:2. For each set every possible pair of consecutive context and prediction windows of length $T_{c}$ and $T_{p}$, respectively, are used. That is, we use rolling (overlapping) windows for both training and testing. [^4] We also fix a known bug (reported in (Das et al., 2023a)) from (Wu et al., 2021) and followup works, where the last $T_{p}$ samples where discarded. We use ADAM optimizer (Kingma & Ba, 2014) together with the original implementation and hyperparameters of each method. Additional experimental details can be found on Appendix C.
5.1 The Shape of ERM
In this section, we provide an empirical analysis of the distribution of errors across time steps when training models with the conventional ERM framework. This aims to uncover phenomena that cannot be inferred from the commonly reported metrics – which are averaged across the window – and to highlight that generalization may hinder the efficacy of loss shaping approaches.
Models For the same dataset and prediction length, different models can have significantly different error patterns, even when having similar MSEs. This can be observed in Figure 3, which portray the test errors of different models trained on the same datasets. This can also be seen in Table 4, which shows that the standard deviation of step-wise MSEs considerably, while their averages (Table 3) are similar.
Prediction Length We observe that the shape of the loss is similar across different prediction lengths in some settings, such as the one depicted in Figure 2. In Appendix C.1 we provide further quantitative evidence of this phenomenon by computing pairwise correlations between (scaled) errors from different prediction lengths.
Figure 2: Step-wise Mean Test Errors for Autoformer on the Weather dataset. Each plot represents a different prediction length window. The shape of the error across the window is similar across prediction lengths.
Generalization In order to assess the impact of the generalization gap on the test error shape, we analyse whether at any given step smaller mean training errors are associated to smaller test errors. We compute for each model, dataset and prediction length, the Spearman correlation between mean stepwise errors computed on the training and test sets. As shown in Table 1, in several setups we find strong correlations, signaling a monotonic relationship between train and test errorsHowever, we also observe a number of weak and even negative correlations. In particular, Informer and the vanilla transformer, present significantly weaker correlations. In these settings, imposing a desired loss shape by modifying the distribution of errors on the training set can be challenging, because it need not affect performance on the test set in a predictable manner.
Autoformer Informer Reformer Transformer
ECL 96 0.60 0.28 0.91 -0.05
192 0.93 0.14 0.67 0.05
336 -0.09 0.31 0.38 0.10
720 0.75 0.05 0.30 0.09
Exchange 96 0.94 0.34 0.91 -0.04
192 0.64 -0.00 -0.36 -0.81
336 0.90 0.07 -0.58 -0.31
720 0.99 0.40 -0.28 -0.42
Weather 96 -0.28 0.80 0.97 -0.33
192 0.98 0.43 0.95 0.08
336 0.95 0.26 0.87 0.26
720 0.82 0.21 0.68 0.10
Table 1: Spearman Correlations between train and test errors for models trained using ERM. Higher means the relationship is closer to being monotonic.
Figure 3: Test MSE computed at each prediction step for different datasets, models and prediction window lengths. The top row shows results for the weather dataset with a predictive window length of 96 steps, and the second row corresponds to the exchange dataset with a predictive window length of 720 steps. Each plot corresponds to a different architecture, and each curve represents a different training algorithm, we include both the ERM baseline and our method using a constant constraint across the window for all models.
5.2 Constraint Level Choice
The most salient design question is the choice of the constraint level $\boldsymbol{\epsilon}_{i}$. In our experimental setting, we begin with the simplest choice of a constant $\boldsymbol{\epsilon}_{i}=\boldsymbol{\epsilon}$ for all timesteps.
The value of $\boldsymbol{\epsilon}$ should depend on the statistical problem at hand, that is, the model class, loss and data distribution. It will also depend on the sample size since, even if the model has enough capacity to fit an extremely tight constraint, it can overfit to training data while being infeasible at test time. On the other hand, if the constraint level is too loose it will not significantly modify the shape of the loss across the window. Lacking prior information or domain knowledge, we propose using the train and validation loss (see Appendix C) of a model trained using ERM in order to gather insights about sensible values of $\boldsymbol{\epsilon}$.
During preliminary experiments, we observed that the median of the ERM validation loss per time step achieved good results. We also conducted an ablation on constraint levels using the upper and lower quantiles of train and validation losses (see Appendix C.1). We report the results on the values of $\boldsymbol{\epsilon}$ with the lower MSE for each combination of model, dataset and prediction length.
In order to quantify the effect of the imposed constraints, we report Mean Constraint Violation (MCV), which we define as the fraction of time-steps that are infeasible, explicitly,
$\mathcal{MCV}(\hat{y},y,\boldsymbol{\epsilon})=\frac{1}{T_{pred}}\sum^{T_{pred}}_{t=1}\mathrm{max}\{0,\ell(\hat{y}_{t},y_{t})-\boldsymbol{\epsilon}\}.$
An effective constrained model should be able to marginally trade-off MSE for a significant reduction in MCV.
5.3 Constrained Loss Shaping Results
Autoformer Informer Reformer Transformer
Ours Ours+R ERM Ours Ours+R ERM Ours Ours+R ERM Ours Ours+R ERM
ECL 96 0.087 0.109 0.091 0.170 0.193 0.176 0.147 0.152 0.149 0.149 0.169 0.152
192 0.026 0.056 0.017 0.173 0.192 0.184 0.166 0.172 0.172 0.063 0.110 0.070
336 0.042 0.080 0.127 0.159 0.174 0.165 0.198 0.162 0.192 0.157 0.168 0.150
720 0.093 0.227 0.075 0.165 0.226 0.173 0.139 0.148 0.139 0.155 0.173 0.151
Exchange 96 0.022 0.006 0.015 0.834 0.872 0.842 0.917 0.873 0.852 0.008 0.003 0.084
192 0.000 0.000 0.791 1.068 1.089 1.071 1.213 1.496 1.209 1.218 1.139 1.200
336 0.000 0.000 2.411 0.000 0.000 0.000 1.423 1.715 1.700 1.323 1.610 1.382
720 0.778 0.382 1.160 0.000 0.000 0.000 1.460 1.506 1.492 0.000 0.000 0.156
Weather 96 0.000 0.000 0.000 0.186 0.230 0.225 0.000 0.019 0.000 0.047 0.030 0.040
192 0.001 0.001 0.002 0.019 0.040 0.045 0.000 0.018 0.129 0.411 0.458 0.485
336 0.018 0.002 0.013 0.411 0.355 0.355 0.010 0.040 0.289 0.209 0.218 0.086
720 0.000 0.000 0.000 0.802 0.949 0.851 0.478 0.631 0.680 0.708 0.856 0.786
Table 2: Test Mean Constraint Violation using constant constraints and vanilla ERM models across different window lengths. Imposing constraints consistently reduces the MCV metric across datasets and window lengths.
Autoformer Informer Reformer Transformer
Ours Ours+R ERM Ours Ours+R ERM Ours Ours+R ERM Ours Ours+R ERM
ECL 96 0.200 0.220 0.204 0.319 0.342 0.325 0.297 0.303 0.299 0.259 0.278 0.262
192 0.226 0.258 0.215 0.341 0.360 0.352 0.326 0.332 0.332 0.274 0.321 0.281
336 0.249 0.287 0.334 0.348 0.371 0.354 0.365 0.329 0.359 0.287 0.297 0.280
720 0.270 0.410 0.252 0.377 0.438 0.385 0.316 0.326 0.316 0.293 0.312 0.289
Exchange 96 0.179 0.163 0.155 0.872 0.910 0.880 1.025 0.981 0.960 0.593 0.556 0.691
192 0.266 0.267 2.052 1.110 1.131 1.113 1.402 1.685 1.398 1.253 1.174 1.235
336 0.446 0.442 3.442 1.088 1.080 1.633 1.632 1.924 1.909 1.362 1.649 1.421
720 1.510 1.205 1.852 1.170 1.166 2.961 2.063 2.114 2.095 1.029 1.015 2.197
Weather 96 0.300 0.314 0.375 0.395 0.441 0.435 0.361 0.380 0.422 0.462 0.435 0.448
192 0.311 0.322 0.303 0.447 0.491 0.515 0.425 0.483 0.703 0.593 0.640 0.667
336 0.374 0.402 0.359 0.656 0.600 0.600 0.589 0.655 1.010 0.901 0.915 0.749
720 0.442 0.447 0.443 1.051 1.198 1.100 0.761 0.914 0.963 0.924 1.072 1.002
Table 3: Test MSE for constrained and ERM models across different window lengths. We observe that the tradeoff between constraint satisfaction and MSE is low, even improving MSE in many instances.
Autoformer Informer Reformer Transformer
Ours Ours+R ERM Ours Ours+R ERM Ours Ours+R ERM Ours Ours+R ERM
ECL 96 0.015 0.013 0.015 0.013 0.018 0.014 0.011 0.008 0.011 0.007 0.008 0.007
192 0.022 0.020 0.022 0.037 0.029 0.041 0.019 0.012 0.020 0.018 0.014 0.018
336 0.019 0.063 0.043 0.018 0.019 0.022 0.023 0.011 0.022 0.022 0.018 0.020
720 0.030 0.156 0.035 0.020 0.040 0.028 0.007 0.007 0.007 0.018 0.018 0.013
Exchange 96 0.084 0.061 0.080 0.274 0.305 0.278 0.121 0.098 0.115 0.162 0.128 0.167
192 0.447 0.389 1.347 0.299 0.331 0.308 0.148 0.196 0.132 0.375 0.374 0.376
336 9.441 0.843 2.132 0.392 0.376 0.387 0.172 0.269 0.193 0.211 0.304 0.218
720 0.226 0.422 1.038 0.915 1.093 0.946 0.280 0.344 0.314 0.624 0.735 0.610
Weather 96 0.029 0.026 0.060 0.141 0.096 0.118 0.107 0.090 0.101 0.121 0.124 0.126
192 0.069 0.040 0.052 0.156 0.127 0.158 0.144 0.132 0.142 0.206 0.242 0.242
336 0.037 0.067 0.068 0.124 0.171 0.136 0.035 0.191 0.141 0.252 0.262 0.224
720 0.081 0.065 0.079 0.251 0.528 0.364 0.038 0.067 0.039 0.273 0.324 0.289
Table 4: Standard deviation of stepwise errors for runs with $\boldsymbol{\epsilon}$ set to the Median of train Error for ERM. “Ours” denotes the constrained approach and “Ours(R)” adds a resilient relaxation with cost $h(\boldsymbol{\zeta})=2\|{\boldsymbol{\zeta}\|}_{2}^{2}$. Imposing constraints tends to reduce the variance across the predictive window. Adding resilience often achieves even less variability, especially in cases where the STD of constrained methods is high.
Our approach effectively affects the distribution of losses across the window as intended in many cases, although it fails to do so in others, as depicted in Figure 3 and Appendix C.1. Nonetheless, we corroborate empirically that it consistently reduces constraint violation (shown in Table 2) on test data across different transformer-based architectures, prediction lengths and datasets.
Furthermore, this does not come at a cost in terms of test MSE (Table 3), which is similar between ERM and our approach in most scenarios. In several settings, our approach even leads to a reduction in MSE. For example, this effect is illustrated in Figure 3, where the Autoformer model trained with constraints attains a smoother error landscape, with most step-wise errors being lower than their ERM counterpart.
The rationale behind imposing an upper bound at each step is reducing the spread of errors across the windows. In Table 4 we show that the standard deviation (computed across the predicted window) is reduced for the vast majority of settings – with the exception of the vanilla transformer, that as already discussed shows poor generalization –. In addition, in Appendix C.1 we provide examples in which the loss landscape in unseen data changes even when imposing looser constraints that are easily satisfied over the training set – i.e., for which the ERM solution is also feasible–. This suggests that irrespective of the level of constraint imposed, using constraints can contribute to enhancing model performance.
Qualitatively, we can distinguish different scenarios in which our approach leads to improvements. On some settings, it can make training errors feasible possibly at the cost of an increase of average training error (e.g., the first row of Figure 4). Conversely, in other scenarios, it fails to obtain a feasible solution and leads to an increase training error, but still improves test performance due to better generalization, like in the second row of Figure 4.
Among the settings in which we are unable to change the test loss as desired, we also distinguish two common failure modes. The first is when our approach fails to impose the constraints on training data, as in the first row of Figure 5, often associated with more challenging settings. Another one is the lack of generalization, depicted in the second row of Figure 5, which corresponds to settings in which our model effectively changes the loss pattern in the training set, but it does not have the same effect on test data.
Figure 4: Examples of success cases. The first row corresponds to the training and testing losses of ERM and Constrained Autoformer on Exchange Rate data, with window length of 720. The second row is Reformer on Exchange Rate with window length of 336. The red lines are the values of $\boldsymbol{\epsilon}$ used during training.
Figure 5: Two instances of failure cases. The first row is the training and testing errors of Autoformer on Weather data with window length of 336. The second row is Informer on ECL data with window length of 720. The red lines are the values of $\boldsymbol{\epsilon}$ used during training.
@@ 5.4 Resilience
Relaxing constraints should allow for a better average performance (objective), in the same way that imposing constraints should only cause a decrease performance when compared to the unconstrained problem. In the previous section, we have shown that – empirically – the latter does not always hold. In the same manner, the Resilient approach did not consistently attain lower test MSEs (Table 3) than the constrained approach, even though constraints were effectively relaxed – by a considerable amount –. Settings in which relaxing constraints is detrimental to performance thus support the claim that loss shaping constraints can also be beneficial in terms of average MSE in some cases. Examples of this behavior can be observed in Figure 3 for the case of Autoformer and Informer.
In terms of constraint violation, which we measure with respect to the fixed constraint levels ($\boldsymbol{\epsilon}$), the resilient approach should be outperformed by its constrained counterpart. However, we find that both approaches perform similarly.
Finally, we observe that the resilient approach can be advantageous in that it reduces standard deviation (Table 4) across the window considerably in many cases. For example, this can be observed for Autoformer (first columns) across several settings.
@@ 6 Conclusion
This paper introduced a time series forecasting constrained learning framework that aims to find the best model in terms of average performance while imposing a user defined upper bound on the loss at each time-step. Given that we observed that the distribution of loss landscapes varies considerably, we explored a Resilient Constrained Learning (Hounie et al., 2023) approach to dynamically adapt individual constraints during training. We analysed the properties of this problem by leveraging recent duality results, and developed practical algorithms to tackle it. We corroborated empirically that our approach can effectively alter the distribution of the loss across the forecast window.
Although we have focused on transformers for long term time series forecasting, motivated by the empirical finding that their loss varies considerably, the loss shaping constrained learning framework can be extended to other settings. This includes a plethora of models, datasets and tasks where losses are distributed across different time steps or grids.

### v2

@@ 5 Experiments
We conduct extensive evaluations comparing constrained and resilient constrained learning against the customary unconstrained training pipeline across eight model architectures and nine popular datasets. For each dataset, we train models with four different predictive window lengths to evaluate the impact of constraints as the forecasting horizons extended. In total, this amounts to 288 different experiment settings.
Explicitly, the datasets are: Electricity Consumption Load (ECL), Weather, Exchange Rate (Lai et al., 2018), Traffic, Electricity Transformer Temperature (ETT) (two hourly datasets and two every 15 minutes) (Zhou et al., 2021), and Influenza Like Illness (ILI). Following the literature, we train with predictive window lengths of 96, 192, 320, and 720 for each dataset, except for Illness, which is trained to predict with lengths of 24, 36, 48, and 60. For a more detailed explanation of the datasets’ contents and source, refer to the Appendix C.1.
We also include a wide variety of time series prediction models comprising seven transformer-based architectures and one non-transformer architecture. Namely, the transformer models are: Reformer (Kitaev et al., 2019), Autoformer (Wu et al., 2021), Informer (Zhou et al., 2021), Pyraformer (Liu et al., 2021), iTransformer (Liu et al., 2024), Nonstationary Transformer (Liu et al., 2022) and a vanilla transformer architecture (Vaswani et al., 2017). The non-transformer model is FiLM (Zhou et al., 2022a).
We follow the same setup, including preprocessing, hyperparameters and implementation, as described in (Kitaev et al., 2019; Wu et al., 2021; Liu et al., 2021; Zhou et al., 2021; Liu et al., 2022; Zhou et al., 2022a). For our method, we only perform a grid search over six values of the constraint level $\epsilon$. The constrained and resilient loss shaping results presented in this section correspond to the best performing constraint level. We provide an additional ablation analysis on one dataset in Appendix D.5.
Data is split into train, validation, and test chronologically with a ratio of 7:1:2. For each data split we extract every pair of consecutive context and prediction windows of length $T_{c}$ and $T_{p}$. That is, we use rolling (overlapping) windows for both training and testing.  [^2] We also fix a known bug (reported in (Das et al., 2023a)) from (Wu et al., 2021) and followup works, where the last $T_{p}$ samples where discarded. Additional experiment details can be found on Appendix C.
5.1 Loss Shaping
In this section, we demonstrate that our approach effectively reduces performance fluctuations across the window while maintaining comparable average performance. We compute the mean squared errors (MSE) at each predicted timestep and report both the average MSE and the standard deviation across the window, referred to as Window STD.
To ease interpretation and facilitate comparisons across different experimental settings, we normalize errors using the unconstrained ERM baseline. For complete result tables with unnormalized MSE values and additional example cases, refer to Appendix D.
Figure 3: Box plots of percentual changes across all experiments. The left column contains plots of MSE across experiments, and the right column is Window STD. We segment experiments by model and prediction length. The $x$ axes of each box plot are sorted by the mean ERM MSE (better models and easier datasets to the left). The full table with the results for each ERM, constrained and resilient setting can be found in Appendix D.
We do not to use constraint violation as a metric because it can be uninformative when the loss landscape is consistently infeasible due to a large generalization gap and constraint levels are constant. In such cases, constraint violation is as informative as MSE, even when the error distribution over time changes.
Due to the large number of experimental setups, we summarize quantitative results in Figure 4, which illustrates the relative differences in MSE and Window STD between ERM and constrained runs. It shows that loss shaping constrained models typically achieve reductions in Window STD while preserving or even improving average MSE.
Figure 4: Distribution of percentual change in test MSE and Window STD across experiment settings when comparing a run of ERM with a constrained run. On average, STD decreases by 3.47% and MSE by 4.47% respectively.
We then analyze how the performance of our method varies across setups, by grouping results by prediction length, model architecture and dataset. First, it varies significantly across models, as shown in the second row of Figure 3. We conjecture that this variability can be attributed to the inductive biases of certain architectures, which makes it harder to impose the desired loss landscape. Second, the relative effect of our method on both MSE and Window STD remains largely unchanged as the prediction window increases, as shown in the third row of box plots in Figure 3. Since both average MSE and Window STD increase with prediction length, a constant relative effect means that the absolute change increases with prediction length. Appendix D.1 shows several concrete examples of this phenomenon.
To qualitatively assess the impact of our method on the error distribution, we include plots of per time-step losses comparing models trained with ERM and our method in Figure 2. In these settings, our approach effectively affects the distribution of losses across the prediction windows. For example, in the Weather dataset (first row), Autoformer and Pyraformer are trading off overall MSE for a flatter error across the window. In other cases, like the Exchange Rate dataset for the Informer and Transformer models, in addition to showing flatter landscapes, the constrained models perform better than their ERM counterparts overall.
Figure 5: Two instances of failure cases. The first row is the training and testing errors of Autoformer on Weather data with window length of 336. The second row is Informer on ECL data with window length of 720. The gray lines are the values of $\boldsymbol{\epsilon}$ used during training.
While we observe that in many cases imposing constraints is beneficial, we also distinguish two failure modes where we are unable to change test loss as desired. The first is when constraints are not feasible at the end of training, as seen in the first row of Figure 5. This is the motivation behind the resilient approach, presented in the next section. A second failure mode is due to the inherent generalization gap, which results in feasible constraints during training, but no effective loss shaping in test data, such as in the second row of Figure 5. We defer to Appendix D.1 for more in-depth comparisons of ERM and constrained runs.
Another advantage of our primal-dual approach is that multipliers indicate the difficulty of satisfying each constraint. Therefore, we can examine the impact of constraints on loss landscapes on a run by analyzing the loss distributions and multipliers in Figure 6. In this experiment, the training loss increases approximately monotonically over the prediction window. The final solution is infeasible, as the MSE constraints are violated beginning at step 341. Consequently, the optimal multipliers are also high, reflecting the difficulty of satisfying the constraints in the latter part of the window. Despite the growth of the multipliers, the model still fails to satisfy the constraints. This infeasibility during training results in ineffective loss shaping during testing.
Figure 6: Training MSE and multipliers of a constrained model, with the ERM curve for reference. The setting is an Autoformer model for the Exchange Rate dataset, with a predictive window length of 720 steps. The gray line is the constraint, with a value of $\sqrt{\epsilon}=0.484$.
@@ 5.2 Resilience
Resilient constrained learning can effectively overcome feasibility issues in training –like those mentioned in the previous section– by relaxing the most difficult constraints. As shown in Figure 7, when comparing hard constraints with resilient constraints using the same setting as the previous section, the resilient approach not only leads to smaller multipliers by relaxing some constraints, but also yields a solution that is more feasible overall.
Figure 7: Training MSE and multipliers of constrained and resilient models. Resilience finds a feasible solution by relaxing the harder constraints. The setting is the same as in Figure  6.
In addition, we find that the solutions achieved by the resilient approach can also have better generalization when compared to hard constraints, as seen in Figure 8. The resilient model reduces the peak around step 150. This illustrates how resilient learning effectively mitigates the failure mode mentioned in the previous section.
As a result of relaxing constraints and solving feasibility issues during training, resilient models are effective in reducing Window STD in a wide variety of settings. This is supported by our empirical results in Tables 5 and 6. For instance, resilient FiLM achieves the lowest Window STD when compared to constrained and ERM-trained models. Models such as Autoformer, Pyraformer and iTransformer also achieve its lowest Window STD when using Resilient constraints.
Furthermore, we find that some settings also achieve its lowest MSE when using resilient constraints. Such is the case of iTransformer on datasets ETTh2,ETTm1 and ETTm2, where the resilient model has the best overall error in almost all predictive window lengths, as shown in the full MSE results, Tables 3 and  4 in Appendix D.
Figure 8: Train and test MSE of constrained and resilient models. By finding a feasible solution in training, generalization improves. The setting is the same as in Figure 6.
Nonstationary Transformer Pyraformer Reformer
ERM Constant Exp ERM Constant Exp ERM Constant Exp
ECL 96 0.178 0.182 0.170 0.287 0.278 0.274 0.299 0.297 0.298
192 0.186 0.191 0.180 0.294 0.292 0.289 0.332 0.326 0.324
336 0.204 0.204 0.196 0.313 0.302 0.295 0.359 0.365 0.364
720 0.233 0.229 0.228 0.310 0.298 0.296 0.316 0.316 0.318
Table 1: Test MSE of ERM, constant-constrained and exponentially-constrained settings in the ECL dataset, for three model architectures
Nonstationary Transformer Pyraformer Reformer
ERM Constant Exp ERM Constant Exp ERM Constant Exp
ECL 96 0.017 0.014 0.016 0.019 0.015 0.015 0.011 0.011 0.012
192 0.026 0.019 0.020 0.022 0.017 0.022 0.020 0.020 0.018
336 0.031 0.024 0.025 0.034 0.023 0.028 0.022 0.023 0.023
720 0.040 0.037 0.038 0.015 0.016 0.016 0.007 0.007 0.007
Table 2: Window STD of ERM, constant-constrained and exponentially-constrained settings in the ECL dataset, for three model architectures.
5.3 Monotonically Increasing Constraints
The results of the previous sections all use $\boldsymbol{\epsilon}_{i}$ set to a constant value derived from the statistics of training or validation ERM errors. As discussed in Section 3, an alternative is to have the constraints take monotonically increasing values, for instance, from a linear or exponential fit. Having increasingly looser constraints is only logical in the common scenario where errors are expected to grow as we predict farther into the future. Note that we call these constraints exponential because they were set from an exponential fit, but they are linear in the optimization variable.
An empirical evaluation on the ECL dataset with three different models shows this is a viable constraint design. Table 1 contains the MSE for each of the three models: in two cases, exponential achieves the lowest MSE across all windows. With respect to Window STD, Table 2 shows that exponential does not always find the most stable shape. However, this is expected, because the monotonically increasing constraint will not necessarily enforce a reduction in variance across the window.
@@ 6 Conclusion
This paper introduced a time series forecasting constrained learning framework that aims to find the best model in terms of average performance while imposing a user-defined upper bound on the loss at each time-step. Given that we observed that the distribution of loss landscapes varies considerably, we explored a resilient constrained learning approach to dynamically adapt individual constraints during training. We analyzed the properties of this problem by leveraging recent duality results and developed practical algorithms to tackle it. We empirically corroborated that our approach can effectively alter the loss distribution across the forecast window.
Although we have focused on transformers for long-term time series forecasting, motivated by the empirical finding that their loss varies considerably, the loss shaping constrained learning framework can be extended to other settings. This includes a plethora of other models, datasets and tasks where losses are distributed across different time steps or grids. Furthermore, in this work, we focused on constraints characterized by a constant value across all timesteps, but other types of constraints, such as exponentially increasing stepwise constraints that we briefly presented, could also be explored in the future.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.